Help needed in calculus ! Thanx

Selve for y when y = S(2x + 1)(x++ x)che. volle+3)(x - 5) de yoſ sinx (640x) *dbe n yordax laxe. Some
Answer #1

Answex] yo f12371) 62?+x) dze Let x²+x=t (2x+1) dx = dt. 22 9. 11 с the_1002) 4.51242 )(2) da y = 5(2x - 2 2 2 2 2 2 3) dx y=1 y. fsinx) (casx) 2 dr let cost = t - Sinx dx=dt y of t² dt. ☺ yos() (1063.3) de let inal = t I dal=dt. yas+? dt. 9. (199te